Solution Preparation
You need to prepare solutions for the reactions you will perform in the laboratory experiment. You will be given a 0.02M stock solution of KMnO4 and a 0.5M stock solution of H2C2O4. The total volume of the solutions should always be 30 mL. You will use DI water to complete the volume as needed. Complete the table below and copy the results into your lab notebook for use during the lab:

Run Volume KMnO4 (mL) Volume H2C2O4 (mL) [KMnO4] (M) [H2C2O4] (M) Water Volume (mL)
1 8.4 16.8
1 4.2 16.8
1 8.4 8.4
Answer #1

Molarity (M) = moles/liter; it changes upon dilution and mixing with other solutions as they are sources of additional dilution.

Total volume = 30 mL = 0.03 L

initial moles KMnO4 = 8.4 mL x 10-3 L x 0.02M = 0.000168 mol ------------ Run 1 & 3

initial moles KMnO4 = 4.2 mL x 10-3 L x 0.02M = 0.000084 mol ------------- Run 2

initial moles H2C2O4 = 8.4 mL x 10-3 L x 0.5M = 0.0042 mol ------------ Run 3

initial moles H2C2O4 = 16.8 mL x 10-3 L x 0.5M = 0.0084 mol ----------- Run 1 & 2

Run 1:

water volume = 30 - [8.4 - 16. 8] = 4.8 mL

[KMnO4] (M) = 0.000168 mol / 0.03 L = 0.0056 M

[H2C2O4] (M = 0.0084 mol / 0.03 L = 0.28 M

Run 2:

water volume = 30 - [4.2 - 16. 8] = 9.0 mL

[KMnO4] (M) = 0.000084 mol / 0.03 L = 0.0028 M

[H2C2O4] (M = 0.0084 mol / 0.03 L = 0.28 M

Run 3:

water volume = 30 - [8.4 - 8.4] = 13.2 mL

[KMnO4] (M) = 0.000168 mol / 0.03 L = 0.0056 M

[H2C2O4] (M = 0.0042 mol / 0.03 L = 0.14 M

Run Volume KMnO4 (mL) Volume H2C2O4 (mL) [KMnO4] (M) [H2C2O4] (M) Water Volume (mL)
1 8.4 16.8 0.0056 0.28 4.8
2 4.2 16.8 0.0028 0.28 9.0
3 8.4 8.4 0.0056 0.14 13.2

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")